What is the force on a 100kg elevator that is falling freely at 9m/s
1
Expert's answer
2021-08-13T10:20:55-0400

I believe the last part should be "falling freely at 9 m/s²."

So, this is the net force that produces a resultant acceleration of 9 meters per second squared. The force if gravity is mg. Therefore, according to Newton's second law:


"-ma=F-mg, \\
F=mg-ma=80\text N."